OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. Choices: A) 1, 2, 3, 4, 5 B) 1, 2, 3, 5, 7 C) 1, 2, 3, 5, 8 D) 1, 3, 5, 7, 9

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

1st term: f(1) easy 2nd term: f(2) easy 3rd term: f(3) = f(2) + f(1) = ? 4th term: f(4) = f(3) + f(2) = ? you get the idea?

OpenStudy (anonymous):

So you have add the f(2)+f(1)= f(3) and so on? @BangkokGarrett

OpenStudy (anonymous):

yes..that's what f(n) = f(n - 1) + f(n - 2), for n > 2 means.

OpenStudy (anonymous):

Once your n is bigger than 2...namely n = 3...you start using that equation. f(1) and f(2) are given.

OpenStudy (anonymous):

so how would I write the 5th term

OpenStudy (anonymous):

Would it be f(5)=f(3)+f(3)?

OpenStudy (anonymous):

f(5) = f(4) + f(3) You have to figure out f(3) first. Then figure out f(4) Then figure out f(5)...the 5th term

OpenStudy (anonymous):

oh okay thanks for all the help

OpenStudy (anonymous):

sure. no problem

OpenStudy (anonymous):

So the answer will be B: 1,2,3,5,7

OpenStudy (anonymous):

No. The first 4 terms are correct. but f(5) is not 7 f(5 = f(4) + f(3) = 5 + 3 = 8

OpenStudy (anonymous):

oh okay thanks